OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (helder_edwin):

you have the line \[ y=-x+6 \] what is the slope?

OpenStudy (anonymous):

-1/0 I think

OpenStudy (helder_edwin):

no! the slope is the coeficient of x!

OpenStudy (anonymous):

the required line is parallel to y=-x so it has the same slope with y-=-x slope =-1 (y-2)/(x-(-4)) = -1 y-2=-(x+4) y-2=-x-4 y=-x-2

OpenStudy (anonymous):

Umm.....The slope is -1 !

OpenStudy (anonymous):

slope is y/x y=-x y/x=-1 so the slope is -1

OpenStudy (helder_edwin):

yes! so you already have m=-1 and the point \[ (x_0,y_0)=(-4,2) \] the equation you are looking for is \[ y-y_0=m(x-x_0) \] so what do you get?

OpenStudy (anonymous):

y-2=x-(-4)

OpenStudy (helder_edwin):

what about m=-1?

OpenStudy (anonymous):

y-2=-1x-(-4)

OpenStudy (helder_edwin):

you have paretheses missing!

OpenStudy (anonymous):

y-2=-1(x-(-4)

OpenStudy (helder_edwin):

that is right \[ y-2=(-1)(x-(-4))=-x-4 \]

OpenStudy (anonymous):

Ok Thanks !

OpenStudy (helder_edwin):

you are welcome remeber! to determine the equation of a line you need (i) two points (from which you can compute the slope), or (ii) the slope and one point